# A Quantile Nelson-Siegel model

### v1

@@ 5 Conclusions
In this study, we introduce a novel time-varying parameters quantile Nelson-Siegel (TVP-QR-NS) model within a Bayesian framework and apply it to an empirical analysis of the US term structure of interest rates. Our in-sample findings reveal an increased dispersion in yield curve factors across quantiles in the post-recession period. Additionally, temporal variations in the distribution of short- and long-term interest rates are observed during the two most recent recessions. Furthermore, evidence indicates significant fluctuations in the interaction between yield curve factors and the macroeconomy across quantiles. Finally, through an out-of-sample forecasting exercise, we demonstrate the superior performance of our proposed approach in tail risk forecasting across all yields compared to its standard conditional mean counterpart.

### v2

@@ 5 Conclusions
This study introduces a novel time-varying parameter quantile Nelson-Siegel (TVP-QR-NS) model, estimated within a Bayesian framework, to analyse the US term structure of interest rates. The out-of-sample forecasting results demonstrate that the TVP-QR-NS model consistently outperforms conventional benchmarks in forecasting tail risks, particularly at longer horizons and for long-term maturities. Importantly, it is the only model consistently included in the model confidence sets for these maturities and horizons, while standard conditional mean models are frequently excluded. The in-sample analysis reveals significant dispersion in yield curve factors across quantiles, especially following recessionary periods. The analysis also uncovers pronounced shifts in the distributions of short- and long-term interest rates—as well as yields at selected maturities—during the Great Recession and the COVID-19 pandemic, highlighting the importance of capturing variation across both time and quantiles. In addition, strong evidence is found of quantile-dependent dynamics in the relationship between yield curve factors and macroeconomic conditions. Overall, these findings highlight the robustness and practical relevance of the TVP-QR-NS framework for risk management and macro-financial policy analysis.
Appendix A Appendix
A.1 Sampling the Decay Parameter
As mentioned above, it has so far been common practice to fix $\lambda$ to be a particular value. On the other hand, we are agnostic and estimate $\lambda(\tau)$ given the data for each quantile. Given a uniform prior for $\lambda(\tau)$, $p(\lambda(\tau))\sim U(a,b)$, the conditional posterior of $(\lambda(\tau)|\bullet)$ is given by
$(\lambda(\tau)|\bullet)\propto p(\mathbf{Y},\mathbf{w}|\bm{\beta})p(\lambda(\tau)),$
where $a<\lambda(\tau)<b$. Since the support of this conditional density is bounded and non-standard, we draw from this conditional density using a Griddy-Gibbs step (Ritter and Tanner, 1992). This is done using the following steps:
  1. 1.
    Construct a grid with grid points $\hat{\lambda(\tau)}_{1},\ldots,\hat{\lambda(\tau)}_{R}$, where $\hat{\lambda(\tau)}_{1}=a$ and $\hat{\lambda(\tau)}_{R}=b$.
  2. 2.
    Compute $F_{i}=\sum_{j=1}^{i}p(\hat{\lambda(\tau)}_{j}|\bullet)$.
  3. 3.
    Generate $U$ from a standard uniform distribution.
  4. 4.
    Find the smallest positive integer $k$ such that $F_{k}\geqslant U$ and return $\lambda(\tau)=\hat{\lambda(\tau)}_{k}$.
A.2 Additional In-Sample Results
Figure 10: Posterior mean estimates of the short-term rate (Level + Slope) for each of the five quantiles (10%, 25%, 50%, 75%, 90%). NBER recession periods in grey shades.